OpenStudy (anonymous):

Please help! A pair of dice is tossed and the sum of the two dice is calculated. find the probability the sum is less than 11.

OpenStudy (anonymous):

I think its 11/12

OpenStudy (anonymous):

if you think about it, out of the 36 possible pairs of numbers that the dice can show, only (6,6) (2 sixes) gets you a sum more than 11. the rest all give you less than 11 (except (6,5) and (5,6) which are 11 exactly)

OpenStudy (anonymous):

so instead of finding Pr(<11), find 1-Pr(>11) (less work) 1- Pr (>11) = 1- Pr(6,6)-Pr(6,5)-Pr(5,6)

OpenStudy (anonymous):

It would be 11/12

OpenStudy (anonymous):

= 1 - (1/36) - (1/36) - (1/36) = 11/12 so yes you got it